# Age Group Sensitivity Analysis of Epidemic Models: Investigating the Impact of Contact Matrix Structure

### v1

@@ Abstract
Understanding the role of different age groups in disease transmission is crucial for designing effective intervention strategies. A key parameter in age-structured epidemic models is the contact matrix, which defines the interaction structure between age groups. However, accurately estimating contact matrices is challenging, as different age groups respond differently to surveys and are accessible through different channels. This variability introduces significant epistemic uncertainty in epidemic models.
In this study, we introduce the Age Group Sensitivity Analysis (AGSA) method, a novel framework for assessing the impact of age-structured contact patterns on epidemic outcomes. Our approach integrates age-stratified epidemic models with Latin Hypercube Sampling (LHS) and the Partial Rank Correlation Coefficient (PRCC) method, enabling a systematic sensitivity analysis of age-specific interactions. Additionally, we propose a new sensitivity aggregation technique that quantifies the contribution of each age group to key epidemic parameters.
By identifying the age groups to which the model is most sensitive, AGSA helps pinpoint those that introduce the greatest epistemic uncertainty. This allows for targeted data collection efforts, focusing surveys and empirical studies on the most influential age groups to improve model accuracy. As a result, AGSA can enhance epidemic forecasting and inform the design of more effective and efficient public health interventions.
@@ 1 Introduction
Mathematical models are extensively employed to understand and predict the spread of infectious diseases within populations. The use of compartmental models is a well-established and frequently applied method in mathematical epidemiology. In this approach, the population is divided into compartments (e.g., susceptible, infectious, recovered) that are significant in terms of the disease, and the transitions between them are modeled by a system of coupled differential equations.
Such models are effective in capturing the dynamics of infectious diseases by simplifying complex interactions within a population. They allow researchers to quantify key epidemiological parameters, such as the basic reproduction number ($\mathcal{R}_{0}$) and the duration of outbreaks, providing insights into the progression and control of epidemics. Moreover, these models can be extended to include additional compartments or factors, such as vaccination, latency periods, or heterogeneous contact patterns, to better reflect real-world scenarios.
@@
In any mathematical model, the certainty of input factors such as parameters is often compromised by natural variations, measurement errors, or outdated measurement techniques [1]. Generally, parameters are derived from experimental data, and in cases where parameter values are not estimated, they are set to plausible values or ranges based on statistical inference and expert judgment [2]. In epidemic modeling, a common and critical question is how altering the parameters affects the system’s temporal dynamics or descriptive quantities such as the final epidemic size or the basic reproduction number.
Sensitivity analysis, which performs Monte Carlo-style simulations to assess the impact of parameter changes, is a widely used approach to address this issue. Sensitivity analysis measures how variations in input parameter values affect the model output [2, 3, 4, 5]. This concept is used in various fields to identify: (i) parameters that need further research to strengthen the knowledge base; (ii) insignificant parameters; (iii) inputs that majorly contribute to output variability; (iv) parameters highly correlated with the output [6]. Sensitivity analysis approaches can be either local or global.
Local methods are straightforward and quick but only examine the effects of local parameter deviations or a selected trajectory in the parameter space. Conversely, global techniques evaluate the entire parameter space along with interactions between parameters to identify all of the system’s critical points [4]. Such techniques include variance-based methods (e.g., the Sobol’ method and Fourier Amplitude Sensitivity Test) [7], global screening methods (e.g., the Morris method) [8], and sampling-based methods (e.g., Monte Carlo filtering; scatter plots [9], Latin hypercube sampling with partial rank correlation coefficient index) [10]. In this present work we focus on the Latin Hypercube Sampling (LHS) combined with Partial Rank Correlation Coefficient (PRCC).
LHS-PRCC sensitivity analysis is an advanced, efficient, and useful statistical technique commonly utilized in uncertainty analysis to explore a model’s parameter space [11]. The LHS/PRCC sensitivity analysis techniques have been applied to deterministic disease transmission models, as seen in [1, 2, 3, 4, 5, 11].
@@
Simple compartment models average out parameters characteristic of individuals, such as infectivity, recovery, or the mortality rate, and consider the population as homogeneous. In reality, however, these parameters significantly depend on several factors like an individual’s age, social behaviour (contact pattern) or health status. Moreover, different age groups interact in distinct ways within a population: the susceptibility to infection, as well as the severity of disease, may vary significantly by age.
@@
Modeling the population in age-stratified compartments offers several advantages. By understanding how different age groups contribute to disease transmission and who is most at risk of severe outcomes, these models can guide decisions on which groups should be prioritized for vaccination to maximize public health benefits. In scenarios where vaccines are limited, age-stratified models are crucial in determining optimal vaccination strategies or planning non-pharmaceutical interventions (NPIs). Policies like targeted lockdowns, age-specific social distancing measures, or selective school closures may be implemented based on model predictions, reducing the overall burden on society while maintaining effectiveness. Thus, we may obtain a more realistic representation of how diseases spread across different segments of the population by stratifying the population into discrete age groups and incorporating age-dependent variations in parameters and age specific contact patterns into the model.
@@
Social contact patterns in different settings are categorized based on the age groups depicted by the contact matrix (CM). Correctly predicting infection dynamics and the effects of control strategies depends on precisely estimating the values of this contact matrix for the different age groups across various settings. Several detailed studies have employed comprehensive quantitative approaches to evaluate social contact patterns in diverse settings and locations [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. These studies have indicated consistent behavioral traits, i.e., assortative mixing; however, the frequency of contact, the extent of intergenerational mixing, and the nature of mixing can vary across age groups and attitudes [24]. The precision of model output simulations over time for the transmission process is hindered by the significant uncertainty in quantifying social contacts among different age groups in various settings.
@@
The aim of this study is to define a sensitivity measure for age-structured disease models in relation to the total number of contacts within each age group. This approach allows us to identify the age groups that have the greatest influence on the key target parameters of the epidemic in question. To achieve this, we use age-stratified epidemic models and perform the LHS/PRCC technique along with our novel sensitivity measure.
@@
The subsequent chapters are organized as follows: the Methods section details the structure of contact matrices and the various elements of the proposed framework, including the main epidemic modeling concepts required for our work, Latin Hypercube Sampling (LHS), Partial Rank Correlation Coefficient (PRCC), the method for determining p-values, and the aggregation approach of PRCC values. In the Demonstrations section, we illustrate the proposed framework for different setups when parameterizing the epidemic model. Finally, we discuss the results of the proposed framework and identify important opportunities to modify and improve the presented methods.
@@ 2 Methods
The application of the AGSA method requires the specification of three inputs. The first input is the contact matrix (see Section 2.1) describing the interaction structure of the population in question. The second input is the age-structured epidemic model, for which we aim to analyze the sensitivity of the contact matrix elements with respect to appropriately defined objective functions. Such objective functions may include the number of infected individuals or those requiring intensive care at the peak of the epidemic wave, or the cumulative number of deaths over the entire course of the epidemic. The third input consists of the additional parameters of the model.
@@
Subsequently, PRCC values are calculated (see Section 2.4 for the contact matrix samples and the corresponding objective function values, from which sensitivity measures are derived. Finally, the obtained sensitivity measures are aggregated (see Section 2.6) by age group, yielding the sensitivity of each age group with respect to the target variables. The flowchart of the complete framework is illustrated in Fig. 1.
Figure 1: Workflow of the AGSA framework. The input parameters include the contact matrix (CM) describing the interaction structure of the population and additional model parameters. Using Latin Hypercube Sampling (LHS), an ensemble of sampled contact matrices is generated and used in conjunction with the epidemic model to calculate target variable values. Partial Rank Correlation Coefficient (PRCC) analysis is performed on these results to derive sensitivity measures, which are then aggregated by age group to determine group-specific sensitivities. Outputs include sensitivity data and visualizations.
The developed framework was built from the ground up in Python using an object-oriented programming approach, ensuring ease of extension and modification to accommodate a wide range of applications. The implementation utilizes core Python libraries, including numpy, scipy, and smt, to efficiently perform the necessary calculations.
@@ 2.1 Social contact matrices
In Sec. Demonstrations of the present paper we use two social contact matrices: in case of the SEIR model, the CM is by Mosson et al. [13], but in case of the Covid model the CM is derived from Prem et al. [12]. In the following we breafly introduce the latter one. The authors obtained the contact structures of populations in 177 geographical locations. The population of one such location is stratified into $n=16$ age groups. Except for the last age group, which is dedicated to the elderly population, each age group spans five years: 0-4, 5-9, 10-14, $\dots$, 75+, with $N_{i}$ representing the number of people in age group $i\in\{0,\dots,15\}$.
Social contacts are categorized into four environments: home, school, work, and other (which includes common social venues such as malls, pubs, public transport, etc.). For each environment, a dedicated contact matrix is constructed using statistical techniques, specifically Markov chain Monte Carlo simulations and convergence diagnosis. Resulting matrix $\tilde{C}^{H}$ for home, $\tilde{C}^{S}$ for school, $\tilde{C}^{W}$ for work, and $\tilde{C}^{O}$ for other settings. The full contact matrix is $\tilde{C}=\tilde{C}^{\text{H}}+\tilde{C}^{\text{S}}+\tilde{C}^{\text{W}}+\tilde{C}^{\text{O}}$. Contact matrices from [12] for Hungary are displayed in Fig. 2. The elements of the contact matrices will be considered as parameters in the sensitivity analysis.
@@
where $\overline{C}_{i,j}=C_{i,j}\cdot N_{i}$, representing the total number of contacts between age groups. This matrix, called the total contact matrix, is symmetric and therefore contains only $136$ unique elements.
It is important to note that in reality every contact related interventions, such as closing schools, maintaining social distance, and working remotely, primarily affect the elements of $C^{S},C^{W},C^{O}$. This means that interactions at home or within the household should be treated separately. We use the following notation: $C^{\Delta}$, which represents the sum of $C^{\text{S}}$, $C^{\text{W}}$, and $C^{\text{O}}$, hence,
@@ 2.2 Age stratified epidemic models and target variables
The AGSA method can be applied to any age-structured epidemic model to analyze the sensitivity of age groups’ impacts on the model’s target variables. These target variables may include any model output, such as the cumulative number of deceased individuals over the course of the epidemic, the peak hospitalization rate, or the ICU bed demand. We demonstrate our method on 2 age stratified epidemic models with different complexity. The most important features of these models are briefly introduced in Section 3.
@@ 2.3 Latin Hypercube Sampling
In the sampling process, the range of each parameter is divided into $\mathcal{S}$ non-overlapping intervals, each with equal probability based on its density function. A single interval is then selected randomly (without repetition), and a random value is drawn from within that interval. This procedure is repeated for all parameters in the analysis. The resulting LHS matrix has rows that represent individual sample vectors $(p^{(i)}_{1},\dots,p^{(i)}_{K})$, with each vector used as input for a separate simulation. The outputs of these simulations are then generated and the values $v^{(i)}$ of the investigated target variable is obtained for every parameter configuration. This is illustrated in Table 1.
@@
Upon symmetrisation, we have to sample the lower triangular elements of the CM. That is 136 independent parameters in case of the CM introduced in Sec. 2.1. To ensure the sampling process is epidemiologically relevant, we make the following assumptions:
@@
    We assume that no intervention can alter home contacts, since a complete lockdown restricts individuals to their households. Thus, we should only sample the changes for ${C}^{\Delta}={C}^{W}+{C}^{S}+{C}^{O}$ (see Eq. 2.4).
@@
    The target parameters for our PRCC analysis include the reproduction numbers $\mathcal{R}_{0}$, the infected peak, the hospitalized peak, the peak of the intensive care unit, and the final death size. Therefore, to assess the effect of variations on outbreaks, we must ensure that our sampling keeps $\mathcal{R}_{0}\geq 1$. Thus, we only consider sampled contact matrices for which the basic reproduction number is greater than 1.
@@
    We aim to examine the overall sensitivity of all contacts characteristic of a given age group, rather than limiting the analysis to specific contacts stored in workplace, school, or other matrices. Thus, it is more practical to sample the reduction ratios of the elements of $C^{\Delta}$ rather than the elements of the contact matrix directly.
To assess A2, first we obtain the basic reproduction number $\mathcal{R}_{0}$ using the next generation matrix (NGM [35]). Then, we introduce parameter $\kappa$ and sample proportions from $[0,\kappa]$. Our objective is to find the value of $\kappa$ so that, $\mathcal{R}_{0}=1$ with the contact matrix
@@
The mapping between the sampled vector $\bar{W}$ and $M_{\text{ratio}}$ can see seen in Appendix LABEL:c_mtx_map. We adjust the matrices by multiplying elementwise the total matrix ${C}^{\Delta}$ by $\mathbbm{1}-M_{\text{ratio}}$, resulting
@@
where $\mathbbm{1}\in\mathbb{R}^{16\times 16}$ represents the matrix $[1]_{i,j=1}^{16}$, $\odot$ indicates element-wise multiplication. In the extreme case, where the maximum change is applied, $M_{\text{ratio}}=\mathbbm{1}$, resulting in ${C}^{\prime}={C}^{\text{H}}$. Using equation 2.5, it is impossible to reduce the contacts below the home contacts.

### v2

# Age Group Sensitivity Analysis in Age Stratified Epidemic Models: Investigating the Impact of Contact Matrix Structure

@@ Abstract
Understanding the role of different age groups in disease transmission is crucial for designing effective intervention strategies. The contact matrix is a key parameter in age-structured epidemic models, which defines the interaction structure between age groups. However, accurately estimating contact matrices is challenging, as different age groups respond differently to surveys and are accessible through various channels. This variability introduces significant epistemic uncertainty in epidemic models.
In this study, we introduce the Age Group Sensitivity Analysis (AGSA) method, a novel framework for assessing the impact of age-structured contact patterns on epidemic outcomes. Our approach integrates age-stratified epidemic models with Latin Hypercube Sampling (LHS) and the Partial Rank Correlation Coefficient (PRCC) method, enabling a systematic sensitivity analysis of age-specific interactions. Additionally, we propose a new sensitivity aggregation technique that quantifies each age group’s contribution to key epidemic parameters.
AGSA helps pinpoint the age groups to which the model is most sensitive by identifying those that introduce the greatest epistemic uncertainty. This allows for targeted data collection efforts, focusing surveys and empirical studies on the most influential age groups to improve model accuracy. As a result, AGSA can enhance epidemic forecasting and inform the design of more effective and efficient public health interventions.
@@ 1 Introduction
Mathematical models are extensively employed to understand and predict the spread of infectious diseases within populations. Compartmental models are a well-established and frequently applied method in mathematical epidemiology. In this approach, the population is divided into compartments (e.g., susceptible, infectious, recovered) that are significant in terms of the disease, and a system of coupled differential equations models the transitions between them.
Such models effectively capture the dynamics of infectious diseases by simplifying complex interactions within a population. They allow researchers to quantify key epidemiological parameters, such as the basic reproduction number ($\mathcal{R}_{0}$) and the duration of outbreaks, providing insights into the progression and control of epidemics. Moreover, these models can be extended to include additional compartments or factors, such as vaccination, latency periods, or heterogeneous contact patterns, to reflect real-world scenarios better.
@@
In any mathematical model, the certainty of input factors such as parameters is often compromised by natural variations, measurement errors, or outdated measurement techniques [1]. Generally, parameters are derived from experimental data. In cases where parameter values are not estimated, they are set to plausible values or ranges based on statistical inference and expert judgment [2]. In epidemic modeling, a common and critical question is how altering the parameters affects the system’s temporal dynamics or descriptive quantities, such as the final epidemic size or the basic reproduction number.
Sensitivity analysis, which performs Monte Carlo-style simulations to assess the impact of parameter changes, is a widely used approach to address this issue. Sensitivity analysis measures how variations in input parameter values affect the model output [2, 3, 4, 5]. This concept is used in various fields to identify (i) parameters that need further research to strengthen the knowledge base, (ii) insignificant parameters, (iii) inputs that majorly contribute to output variability, (iv) parameters highly correlated with the output [6]. Sensitivity analysis approaches can be either local or global.
Local methods are straightforward and quick but only examine the effects of local parameter deviations or a selected trajectory in the parameter space. Conversely, global techniques evaluate the entire parameter space along with interactions between parameters to identify all of the system’s critical points [4]. Such techniques include variance-based methods (e.g., the Sobol’ method and Fourier Amplitude Sensitivity Test) [7], global screening methods (e.g., the Morris method) [8], and sampling-based methods (e.g., Monte Carlo filtering; scatter plots [9], Latin hypercube sampling with partial rank correlation coefficient index) [10]. This work focuses on the Latin Hypercube Sampling (LHS) combined with the Partial Rank Correlation Coefficient (PRCC).
LHS-PRCC sensitivity analysis is an advanced, efficient, and useful statistical technique commonly utilized in uncertainty analysis to explore a model’s parameter space [11]. The LHS/PRCC sensitivity analysis techniques have been applied to deterministic disease transmission models, as seen in [1, 2, 3, 4, 5].
@@
Simple compartment models average out parameters characteristic of individuals, such as infectivity, recovery, or mortality rate, and consider the population as homogeneous. However, these parameters significantly depend on several factors, such as an individual’s age, social behavior (contact pattern), or health status. Moreover, different age groups interact in distinct ways within a population: the susceptibility to infection and the severity of disease may vary significantly by age.
@@
Modeling the population in age-stratified compartments offers several advantages. By understanding how different age groups contribute to disease transmission and who is most at risk of severe outcomes, these models can guide decisions on which groups should be prioritized for vaccination to maximize public health benefits. In scenarios where vaccines are limited, age-stratified models are crucial in determining optimal vaccination strategies or planning non-pharmaceutical interventions (NPIs). Policies like targeted lockdowns, age-specific social distancing measures, or selective school closures may be implemented based on model predictions, reducing the overall burden on society while maintaining effectiveness. Thus, we may obtain a more realistic representation of how diseases spread across different population segments by stratifying the population into discrete age groups and incorporating age-dependent variations in parameters and age-specific contact patterns into the model.
@@
Social contact patterns in different settings are categorized based on the age groups depicted by the contact matrix (CM). Correctly predicting infection dynamics and the effects of control strategies depends on precisely estimating the values of this contact matrix for the different age groups across various settings. Several detailed studies have employed comprehensive quantitative approaches to evaluate social contact patterns in diverse settings and locations [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. These studies have indicated consistent behavioral traits, i.e., assortative mixing; however, the frequency of contact, the extent of intergenerational mixing, and the nature of mixing can vary across age groups and attitudes [25]. Significant uncertainty hinders the precision of model output simulations over time for the transmission process, quantifying social contacts among different age groups in various settings.
@@
This study aims to define a sensitivity measure for age-structured disease models in relation to the total number of contacts within each age group. This approach allows us to identify the age groups that have the most significant influence on the epidemic’s key target parameters. We use age-stratified epidemic models to achieve this and perform the LHS/PRCC technique and our novel sensitivity measure.
@@
The subsequent chapters are organized as follows: the Methods section details the structure of contact matrices and the various elements of the proposed framework, including the main epidemic modeling concepts required for our work, Latin Hypercube Sampling (LHS), Partial Rank Correlation Coefficient (PRCC), the method for determining p-values, and the aggregation approach of PRCC values. In the Demonstrations section, we illustrate the proposed framework for different setups when parameterizing the epidemic model. Finally, we discuss the proposed framework’s results and identify essential opportunities to modify and improve the presented methods.
@@ 2 Methods
The application of the AGSA method requires the specification of three inputs. The first input is the contact matrix (see Section 2.1) describing the interaction structure of the population in question. The second input is the age-structured epidemic model, for which we aim to analyze the sensitivity of the contact matrix elements concerning appropriately defined objective functions. Such objective functions may include the number of infected individuals, those requiring intensive care at the peak of the epidemic wave, or the cumulative number of deaths over the entire epidemic. The third input consists of the additional parameters of the model.
@@
Subsequently, PRCC values are calculated (see Section 2.4 for the contact matrix samples and the corresponding objective function values, from which sensitivity measures are derived. Finally, the obtained sensitivity measures are aggregated (see Section 2.6) by age group, yielding the sensitivity of each age group concerning the target variables. The flowchart of the complete framework is illustrated in Fig. 1.
Figure 1: Workflow of the AGSA framework. The input parameters include the contact matrix (CM) describing the interaction structure of the population and additional model parameters. Using Latin Hypercube Sampling (LHS), an ensemble of sampled contact matrices is generated and used with the epidemic model to calculate target variable values. Partial Rank Correlation Coefficient (PRCC) analysis is performed on these results to derive sensitivity measures, which are then aggregated by age group to determine group-specific sensitivities. Outputs include sensitivity data and visualizations.
The framework was developed from the ground up in Python using an object-oriented programming approach. This ensures ease of extension and modification to accommodate a wide range of applications. The implementation utilizes core Python libraries, including numpy, scipy, and smt, to perform the necessary calculations efficiently.
@@ 2.1 Social contact matrices
In Sec. Demonstrations of the present paper we use two social contact matrices: in the case of the SEIR model, the CM is by Mosson et al. [13], but in the case of the Covid model, the CM is derived from Prem et al. [12]. In the following, we briefly introduce the latter one. The authors obtained the contact structures of populations in 177 geographical locations. The population of one such location is stratified into $n=16$ age groups. Except for the last age group, which is dedicated to the elderly population, each age group spans five years: 0-4, 5-9, 10-14, $\dots$, 75+, with $N_{i}$ representing the number of people in age group $i\in\{0,\dots,15\}$.
Social contacts are categorized into four environments: home, school, work, and other (which includes typical social venues such as malls, pubs, public transport, etc.). A dedicated contact matrix for each environment is constructed using statistical techniques, specifically Markov chain Monte Carlo simulations and convergence diagnosis. Resulting matrix $\tilde{C}^{H}$ for home, $\tilde{C}^{S}$ for school, $\tilde{C}^{W}$ for work, and $\tilde{C}^{O}$ for other settings. The full contact matrix is $\tilde{C}=\tilde{C}^{\text{H}}+\tilde{C}^{\text{S}}+\tilde{C}^{\text{W}}+\tilde{C}^{\text{O}}$. Contact matrices from [12] for Hungary are displayed in Fig. 2. The elements of the contact matrices will be considered as parameters in the sensitivity analysis.
@@
where $\overline{C}_{i,j}=C_{i,j}\cdot N_{i}$, representing the total number of contacts between age groups. This matrix, the total contact matrix, is symmetric and contains only $136$ unique elements.
It is important to note that in reality, every contact-related intervention, such as closing schools, maintaining social distance, and working remotely, primarily affects the elements of $C^{S},C^{W},C^{O}$. This means that interactions at home or within the household should be treated separately. We use the following notation: $C^{\Delta}$, which represents the sum of $C^{\text{S}}$, $C^{\text{W}}$, and $C^{\text{O}}$, hence,
@@ 2.2 Age stratified epidemic models and target variables
The AGSA method can be applied to any age-structured epidemic model to analyze the sensitivity of age groups’ impacts on the model’s target variables. These target variables may include any model output, such as the cumulative number of deceased individuals throughout the epidemic, the peak hospitalization rate, or the ICU bed demand. We demonstrate our method on two age-stratified epidemic models with different complexity. The most important features of these models are briefly introduced in Section 3.
A brief note on using the basic reproduction number as a target variable. To determine the model’s base transmission rate, we assume the $\mathcal{R}_{0}$ value the disease would have in the absence of interventions. Given this predefined $\mathcal{R}_{0}$, along with the model parameters and the contact matrix, the base transmission rate $\beta_{0}$ can be calculated. In real-life situations, contact-based interventions do not change this $\beta_{0}$ as it is a parameter characteristic of the disease.
During sensitivity analysis, modifying the contact matrix elements alters $\mathcal{R}_{0}$ since the next-generation matrix (see Sec. 2.3) computation incorporates these elements. As a result, $\mathcal{R}_{0}$ can also be selected as a target variable, allowing us to assess how changes in contact patterns influence the basic reproduction number.
@@ 2.3 Latin Hypercube Sampling
In the sampling process, the range of each parameter is divided into $\mathcal{S}$ non-overlapping intervals, each with equal probability based on its density function. Then, a single interval is selected randomly (without repetition), and a random value is drawn from within that interval. This procedure is repeated for all parameters in the analysis. The resulting LHS matrix has rows that represent individual sample vectors $(p^{(i)}_{1},\dots,p^{(i)}_{K})$, with each vector used as input for a separate simulation. The outputs of these simulations are then generated, and the values $v^{(i)}$ of the investigated target variable are obtained for every parameter configuration. This is illustrated in Table 1.
@@
Upon symmetrization, we have to sample the lower triangular elements of the CM. That is 136 independent parameters in the case of the CM introduced in Sec. 2.1. To ensure the sampling process is epidemiologically relevant, we make the following assumptions:
@@
    We assume that no intervention can alter home contacts since a complete lockdown restricts individuals to their households. Thus, we should only sample the changes for ${C}^{\Delta}={C}^{W}+{C}^{S}+{C}^{O}$ (see Eq. 2.4).
@@
    The target parameters for our PRCC analysis include the reproduction numbers $\mathcal{R}_{0}$, the infected peak, the hospitalized peak, the peak of the intensive care unit, and the final death size. Therefore, to assess the effect of variations on outbreaks, we must ensure that our sampling keeps $\mathcal{R}_{0}\geq 1$. Thus, we only consider sampled contact matrices for which the basic reproduction number exceeds 1.
@@
    We aim to examine the overall sensitivity of all contacts characteristic of a given age group rather than limiting the analysis to specific contacts stored in the workplace, school, or other matrices. Thus, it is more practical to sample the reduction ratios of the elements of $C^{\Delta}$ rather than the elements of the contact matrix directly.
To assess A2, first, we obtain the basic reproduction number $\mathcal{R}_{0}$ using the next generation matrix (NGM [36]). Then, we introduce parameter $\kappa$ and sample proportions from $[0,\kappa]$. Our objective is to find the value of $\kappa$ so that, $\mathcal{R}_{0}=1$ with the contact matrix
@@
The mapping between the sampled vector $\bar{W}$ and $M_{\text{ratio}}$ can be seen in Appendix C. We adjust the matrices by multiplying elementwise the total matrix ${C}^{\Delta}$ by $\mathbbm{1}-M_{\text{ratio}}$, resulting
@@
where $\mathbbm{1}\in\mathbb{R}^{16\times 16}$ represents the matrix $[1]_{i,j=1}^{16}$, $\odot$ indicates element-wise multiplication. In the extreme case, where the maximum change is applied, $M_{\text{ratio}}=\mathbbm{1}$, resulting in ${C}^{\prime}={C}^{\text{H}}$. Using equation 2.5, reducing the contacts below the home contacts is impossible.